OpenStudy (anonymous):

Solve m plus one-fourth equals three-fourths. m equals negative one-half m equals one-fourth m equals one-half m equals one

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

m + 1/4 = 3/4 (subtract 1/4 from each side) m = 3/4 - 1/4 m = 2/4 = 1/2